OpenStudy (anonymous):

sugahead12 Of the 1,000 students in a local college, 420 own brand X mobile phones and 580 own brand Y mobile phones. Of these students, 80 own both brands of mobile phones. Find the probability that a student chosen at random has a brand X mobile phone given that he has a brand Y mobile phone. A. 2/14 B. 5/21 C. 3/28 D. 4/29

OpenStudy (anonymous):

@sugahead12

OpenStudy (anonymous):

He dont know

OpenStudy (anonymous):

@princeharryyy

OpenStudy (anonymous):

@OregonDuck

OpenStudy (anonymous):

@dmart180

OpenStudy (mathmate):

Have you learned conditional probability?

OpenStudy (anonymous):

Its confusing

OpenStudy (mathmate):

@Jade3115 Do you know how to draw the Venn diagram of the situation?

OpenStudy (mathmate):

 To draw the Venn diagram, you subtract 80 (own both) from 420 who own phone X to get 340 students who own ONLY phone x. Similarly for y. Let X=event that a student owns phone X, and Y=event that a student owns phone Y. So out of the 1000 students, we have 80 who own both, so P(X\(\cap\)Y)=80/1000. P(Y)=580/1000. And so conditional probability that the student owns X given that he owns Y is, by definition of conditional probability P(X|Y)=P(X\(\cap\)Y)/P(Y) Since all quantities on the right-hand side are known, the required probability can be calculated.

OpenStudy (anonymous):

Thanks so much ! Can you help me with another ?

OpenStudy (mathmate):

Definitely, if you participate!

OpenStudy (anonymous):

in a factory the chance that a certain machine works without overheating in the morning is 50% . if it runs smoothly all morning then there is an 85% chance that it will continue for the rest of the day and 15% chance that it will stop due to over heating . what is the probability that on a given gay it will work in the morning and overheat later on ?

OpenStudy (anonymous):

& my bad I'm doing this Plato thing

OpenStudy (mathmate):

 Here's a tree diagram that I started. Can you please fill in the probabilities of each branch?

OpenStudy (anonymous):

Its 7.5% ?

OpenStudy (mathmate):

Did you get that without using the tree diagram?

OpenStudy (anonymous):

I did sum extra stuff & it seemed right . & it was correct !! Any idea ab this one ? Two friends, Bob and Ben, each buy one lottery ticket. Each ticket contains six numbers from a total of one hundred numbers (0–99). Bob chooses the numbers 1, 2, 3, 4, 5, 6, and Ben chooses the numbers 39, 45, 66, 72, 74, 89. Who has a higher probability of winning?

OpenStudy (mathmate):

What do you think?

OpenStudy (anonymous):

I think Ben just not sure

OpenStudy (mathmate):

and your reasoning?

OpenStudy (anonymous):

Bc his numbers are kinda scattered so he has a better number choices

OpenStudy (mathmate):

You know in these lotteries, they use identical physical balls each with a number painted on it. This is to make sure every number has an equal chance of being drawn. So based on this information, can you reconsider your reasoning?

OpenStudy (anonymous):

Not enough information ?

OpenStudy (anonymous):

Noo both have equal chances right ?

OpenStudy (anonymous):

Yes it was lol thank you